A garden's length is five times its width. Write an expression for the number of sections of 10 foot fence that are needed for this garden.
Answer by RAY100(1637)

Perimeter = 2* Length + 2* Width
BUT length = 5 width
P = 2 * (5w) + 2w = 10 w + 2 w = 12 w
Number of lengths = (12 *w )/10 = 1.2 w
( need more for actual number)